The perimeter P of a rectangle with length L and width w is given by the equation P=2(L+w). Which of the equation gives the width in terms of P and L?
A. w= P-2L
B. w=2(P-2L)
C. w=(P-2L)/2 >>>fraction...parentheses are over two.
D. w= (P+2L)/2 >>>same^

Answer by jim_thompson5910(35256)

P=2(L+w)

P=2L+2w

P-2L=2w

2w = P-2L

w = (P-2L)/2

So the answer is choice C